OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (anonymous):

@ganeshie8 @UnkleRhaukus

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

@jod1n3lson

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

-8-16+6+8+8

OpenStudy (anonymous):

-24+21

OpenStudy (anonymous):

the answer is -3

OpenStudy (anonymous):

got it?

OpenStudy (haseeb96):

the answer is 9 because after putting values u got 32 -32 + 9 -16 +16 so 32 and 16 cancel out then 9 will be remaining

OpenStudy (unklerhaukus):

\[2x^2 - 2z^4 + y^2 - x^2 + z^4\Big|\quad x = -4,\quad y = 3,\quad z = 2\\=2(-4)^2 - 2(2)^4 +(3)^2 - (-4)^2 + (2)^4\\= 2\times16-2\times16+9-16+16\\=\dots\]